OpenStudy (anonymous):

The temperature of a substance in a laboratory beaker dropped thirteen and one-half degrees Fahrenheit from its starting temperature, ending at sixty-eight and three-fourths degrees Fahrenheit. Let s equal the starting temperature. Which mathematical equation could be used to determine the starting temperature of the substance? s plus thirteen and one-half equals sixty-eight and three-fourths s plus sixty-eight and three-fourths equals thirteen and one-half s minus thirteen and one-half equals sixty-eight and three-fourths s times sixty-eight and three-fourths equals thirteen and o

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

The start temperature is s, and the temperature dropped thirteen and one-half degrees Fahrenheit. So first we get \[13\frac{ 1 }{ 2 }\] The ending temperature is sixty-eight and three-fourths degrees Fahrenheit. Thus \[s-13\frac{ 1 }{ 2 }=68\frac{ 3 }{ 4 }\] is the equation about s.

OpenStudy (anonymous):

@iceicebaby

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

: )